# Learning General-Purpose Biomedical Volume Representations using Randomized Synthesis

### v1

@@ Abstract
Current volumetric biomedical foundation models struggle to generalize as public 3D datasets are small and do not cover the broad diversity of medical procedures, conditions, anatomical regions, and imaging protocols. We address this by creating a representation learning method that instead anticipates strong domain shifts at training time itself. We first propose a data engine that synthesizes highly variable training samples that enable generalization to new biomedical contexts. To then train a single 3D network for any voxel-level task, we develop a contrastive learning method that pretrains the network to be stable against nuisance imaging variation simulated by the data engine, a key inductive bias for generalization. This network’s features can be used as robust representations of input images for downstream tasks and its weights provide a strong, dataset-agnostic initialization for finetuning on new datasets. As a result, we set new standards across both multimodality registration and few-shot segmentation, a first for any 3D biomedical vision model, all without (pre-)training on any existing dataset of real images.
@@ 1 Introduction
Biomedical vision models trained on imaging studies with fixed protocols rarely generalize to new populations, medical procedures, and imaging devices. These domain shifts then necessitate clinically infeasible reannotation and retraining cycles, especially for adaptation to new tasks. Further, volumetric annotated biomedical datasets are especially limited in sample size and focused on specific medical procedures, diseases, or scales of anatomy, leading current networks to overfit to a small subset of biomedical tasks. To overcome this data scarcity and heterogeneity, we present a representation learning framework driven by a synthetic data engine. Our approach yields a generalist 3D network that performs well on diverse voxel-level tasks across a range of unseen biomedical contexts.
@@
Current biomedical foundation models are trained by aggregating publicly available datasets to cover multiple domains (Butoi et al., 2023; Chen et al., 2024a; Liu et al., 2023a; Ma & Wang, 2023; MH Nguyen et al., 2024; Pachitariu & Stringer, 2022; Xie et al., 2022). However, persistent gaps hinder their widespread adoption. For example, some methods operate only on specific modalities and regions that can be tractably scaled up in sample size, such as chest X-ray (Chen et al., 2024b), and thus cannot learn general representations for most other domains, such as in utero fetal MRI. Others treat 2D slices of volumetric images as independent data points (Butoi et al., 2023; Ma & Wang, 2023), often constructing training sets with high inter-sample correlation yielding models that fail to produce consistent 3D results. Further, existing foundation models almost exclusively focus on segmentation and classification and neglect other key vision tasks such as registration. To our knowledge, no biomedical vision foundation model has been demonstrated for multiple disparate 3D tasks yet.
@@
Contributions. This paper makes advances on two fronts. To gain robustness to large domain shifts in downstream deployment, we first propose a biomedically informed data engine whose samples encompass a wide range of appearances and semantics. This engine uses randomly sampled spatial configurations of biomedical shape templates to synthesize images with arbitrary resolutions, appearances, imaging physics, and crucially, minimal influence from any existing biomedical dataset. Unlike training on samples from GANs or diffusion models, which are limited to reproducing only their original training distribution, our engine synthesizes highly diverse samples useful for arbitrarily new biomedical contexts that we do not have training data for. We then develop a contrastive learning framework that uses paired samples from the data engine to pretrain a network for general voxel-level tasks using an inductive bias of approximate stability to nuisance imaging variation that does not change image semantics, a key property for generalization across datasets (Gruver et al., 2022).
@@
We experimentally demonstrate that the resulting features and weights enable broad generalization on the key biomedical tasks of 3D registration and segmentation across several diverse datasets. We achieve state-of-the-art unsupervised multimodality image registration by simply using the network’s approximately appearance invariant and pose equivariant representations (Fig. 1) to drive existing registration solvers. The proposed network can also be used as an off-the-shelf dataset-agnostic initialization for finetuning on any voxel-level task. Specifically, we demonstrate strong few-shot segmentation performance in a few-shot setting across highly diverse downstream datasets, thereby removing the need for cumbersome dataset-specific self-supervised pretraining. Our code and trained model weights are available at https://github.com/neel-dey/anatomix.
@@ 2 Related work
Figure 2: Data engine. A. We randomly sample binary labels as templates from a large database of segmentations to create 3D shape ensembles of randomly deformed templates. B. Given a synthesized shape ensemble and an appearance model, we synthesize two volumes to pretrain a network with a dense contrastive objective. C. Example synthetic training volumes.
@@
Domain randomization. To learn robustness to domain shifts at deployment, domain randomized generative models (Tobin et al., 2017) trade realism for diversity when generating training data for downstream models. For example, domain randomized methods for brain segmentation (Billot et al., 2023a; b; Gopinath et al., 2023; Hoopes et al., 2022b) and registration (Hoffmann et al., 2021; Iglesias, 2023) train on synthetic brains simulated from label maps and require large collections of expert brain annotations. Recent work in dataset-agnostic instance segmentation (Chollet et al., 2024; Dey et al., 2024) generalizes beyond brains by simulating both annotations and images using a pre-specified shape prior. We build on these concepts to train a task-agnostic network to simulate images with highly variable appearances and physics from compositions of biomedical shape templates.
@@
Volumetric pretraining. Many methods pretrain on unannotated images from a dataset before supervised finetuning on a small labeled subset. These methods often employ self-supervised reconstructive (Chen et al., 2019a; Tang et al., 2022; Valanarasu et al., 2024; Zhou et al., 2023; 2021) and/or discriminative (Chaitanya et al., 2020; 2023; Ren et al., 2022; You et al., 2024) losses. However, these pretraining strategies exploit heuristics about their target datasets that often do not broadly generalize (Dong et al., 2021; Ren et al., 2022). Our approach instead yields a network that generalizes to arbitrary datasets and does not require bespoke pretraining frameworks for each project. Lastly, recent biomedical foundation models trained on pooled datasets of 2D slices (Butoi et al., 2023; Ma & Wang, 2023; MH Nguyen et al., 2024; Wong et al., 2023) generally require interaction (via bounding boxes, scribbles, etc.), struggle with 3D consistency, and are restricted to segmentation. We instead directly train for general tasks using synthetic 3D volumes and do not work on interactive tasks.
@@ 3 Methods
Figure 3: Representation learning. Given a 3D label map and two synthetic volumes generated from it, we process them with a single UNet (with shared weights). The UNet is pretrained contrastively at each layer of the decoder. For a randomly sampled anchor, features sampled from corresponding labels in both volumes are treated as positives and features from other labels are considered negatives.
@@
Data engine: label ensemble model (Fig. 2A). We create synthetic 3D label ensemble volumes by sampling from a repository of biomedical shape templates. As templates, we use the freely available $\sim$45,000 binary volumes from the TotalSegmentator dataset of 104 annotated organs in 1,204 CT volumes (Wasserthal et al., 2023). For each label ensemble volume, we iteratively populate a 3D volume with a random number of randomly sampled templates that are each then deformed and assigned the label corresponding to the sampling iteration. To simulate anatomy being surrounded by empty space (as is common in radiology), we apply a foreground mask to two-thirds of the synthesized label ensembles by multiplying them with a randomly deformed binary sphere with a random radius and center. Finally, we randomly encase half of the foreground-masked volumes within envelope labels of random widths, to emulate layer-like structures common to some biomedical applications (e.g., fat).
Data engine: appearance model (Fig. 2B). Given a 3D label ensemble $L$ with $K$ labels, we sample the intensities of two volumes $V_{1}$ and $V_{2}$ from two independent $K$-component Gaussian mixture models (GMMs) each with parameters $\{\mu_{k1},\sigma_{k1}^{2}\}_{k=1}^{K}$ and $\{\mu_{i2},\sigma_{i2}^{2}\}_{k=1}^{K}$, respectively, all of which are randomly drawn from uniform distributions. For each spatial index in $L$ with label $k$, we sample the initial intensities in $V_{1}$ and $V_{2}$ from $\mathcal{N}(\mu_{k1},\sigma_{k1}^{2})$ and $\mathcal{N}(\mu_{k2},\sigma_{k2}^{2})$, respectively. We then pointwise multiply them with Perlin noise (Perlin, 1985) to simulate spatial texture and augment using transformations relevant to biomedical volumes such as random bias fields, Fourier-spikes, Gamma shifts, blurring, Gibbs ringing, resolution degradations, noise, motion, flips, and affine warps. All intensity augmentations are sampled independently but the geometric augmentations are shared.
@@
Contrastive pretraining (Fig. 3). As our data engine provides exact label supervision, we develop a representation learning loss that is a spatial extension of multi-positive supervised contrastive learning (Khosla et al., 2020). To pretrain network $F:\mathbb{R}^{H\times W\times D}\rightarrow\mathbb{R}^{H\times W\times D\times C}$ where $H,W,$ and $D$ are spatial dimensions and $C$ is the number of output features, we use an inductive bias of voxels within a 3D shape having similar spatial representations, regardless of appearance. We assume that an anchor spatial index $i\in I$ (where $I=\{1,\dots,2HWD\}$, i.e., voxels pooled from $V_{1}$ and $V_{2}$) with features $f_{i}\in\mathbb{R}^{C}$ in label $k$ should have similar representations to all other indices in label $k$ in both $F(V_{1})$ and $F(V_{2})$ and dissimilar representations to indices from other labels. As in (Chen et al., 2020a), we use a non-linear projection $Z:\mathbb{R}^{H\times W\times D\times C}\rightarrow\mathbb{R}^{H\times W\times D\times C_{Z}}$ on $F$’s outputs followed by an $L_{2}$-normalization when computing the contrastive loss,
@@
Figure 4: Multi-modality 3D registration. The ConvexAdam registration solver (Siebert et al., 2021) driven by our network features (“Ours”) accurately aligns challenging intra-subject abdominal MRI-CT (top) and inter-subject cardiac MRI-CT (bottom) 3D multi-modality pairs with large deformations.
@@
Downstream tasks: multi-modality registration. Gradient-based deformable registration objectives typically take the form of $\mathcal{L}_{reg}=d(V_{\mathrm{fixed}},V_{\mathrm{moving}}\circ\varphi)+\lambda\mathrm{Reg}(\varphi)$, where image $V_{\mathrm{moving}}$ is to be aligned to $V_{\mathrm{fixed}}$ by deformation $\varphi$ (subject to regularization $\mathrm{Reg}(\cdot)$) and $d(\cdot)$ is an image dissimilarity score. To align images across imaging modalities, we simply replace the input volumes with our pretrained network’s features as in $\mathcal{L}_{reg}=d(F(V_{\mathrm{fixed}}),F(V_{\mathrm{moving}})\circ\varphi)+\lambda\mathrm{Reg}(\varphi)$. This approach is compatible with any existing high-performance registration solver, such as the ConvexAdam (Siebert et al., 2021) and ANTs (Tustison et al., 2020) frameworks used in our experiments below.
Downstream tasks: few-shot segmentation. For $N$-label segmentation, we use a small set of annotated volumes to finetune the pretrained network $F$, with an additional convolutional layer with softmax activation with $N$ output channels. We optimize the network using an equally weighted sum of the soft Dice and cross-entropy losses (Isensee et al., 2021; Taghanaki et al., 2019). To extract strong performance for all baselines in the challenging setting of finetuning on only one or few annotated volumes, we use extensively-tuned augmentation pipelines and finetune all layers of each network for a high number of iterations (37,500) with cosine learning rate decay.
@@ 4 Experiments
The pretrained network’s output features for inter and intra-subject volume pairs with semantically similar content are visualized in Fig. 1. Below, we present experiments that investigate the utility of our learned representations for multi-modality registration, the network weights as a pretrained initialization for few-shot segmentation, and analyze and ablate our framework’s modeling decisions.
@@ 4.1 Unsupervised multi-modality deformable 3D registration
Data and setup. We use the Learn2Reg AbdomenMRCT (Hering et al., 2021) (L2RAb) and MM-WHS (Gao et al., 2023; Zhuang, 2018; Zhuang et al., 2019) datasets to benchmark MRI to CT volume registration. L2RAb is an abdominal registration benchmarking dataset, whose publicly available portion provides eight affine-aligned intra-subject MRI and CT pairs of size $192\times 160\times 192$ at $2\times 2\times 2\mathrm{mm}^{3}$ resolution, with labels for four organs. MM-WHS, originally a heart segmentation dataset, contains 20 annotated MRIs and CTs (from distinct subjects) and we affine-align all volumes to a common space of size $160\times 160\times 128$ at $1.142\times 1.142\times 1.283\mathrm{mm}^{3}$ resolution (see App. B.4.4 for further details). The registration experiments are unsupervised and we therefore split L2RAb and MM-WHS into 1/7 and 5/15 validation/testing pairs, respectively, where the validation pair(s) are only used for tuning registration hyperparameters.
Baselines and evaluation. We compare unsupervised and training-free multimodality registration frameworks. Our iterative baselines include the widely-used ANTs library (Avants et al., 2008; Tustison et al., 2020) with mutual information loss (Mattes et al., 2001; Wells III et al., 1996) (ANTs-MI) and the state-of-the-art (Hering et al., 2021) multimodality method, ConvexAdam (Siebert et al., 2021). We further use two dataset-agnostic registration networks, SynthMorph-shapes (Hoffmann et al., 2021) and uniGradICON (Tian et al., 2024a), with the latter also using optional instance optimization (uniGradICON+IO). For evaluation, we report post-registration volume overlap (Dice) of anatomical structures. We also assess deformation inverse consistency using the percentage of folding voxels, where $\mathrm{det}(J_{\varphi})<0$, for Jacobian $J_{\varphi}$ of the estimated $\varphi$. Folding percentages below 0.5% of all voxels are generally considered negligible and methods producing higher Dice values while staying under this threshold are preferred (Dey et al., 2022; Qiu et al., 2021).
@@
Results. Figs. 4 and 5 present results on held-out testing pairs. ANTS-Ours strongly improves upon the typically-used ANTS-MI, with 26 and 5 points of median Dice improvement on L2RAb and MM-WHS, respectively, while maintaining nearly identical low folding characteristics. Further, driven by our network features, ConvexAdam-Ours outperforms all methods in terms of volume overlap and improves on its base ConvexAdam method by 11 and 6 Dice points, with the same folding ratios.
In contrast, the SynthMorph-shapes and uniGradICON+IO methods perform well on MM-WHS (where all hearts are roughly centered) but cannot handle the larger deformations in L2RAb. They do, however, yield nearly diffeomorphic transformations, producing almost zero folds. Lastly, without iterative optimization, uniGradICON demonstrates limited generalization across large intensity-based domain gaps. We note that all methods produce folding percentages that are under the threshold of 0.5% folding voxels. Additional grid search results on the validation sets are in App. A.2.

### v2

@@ Abstract
Current volumetric biomedical foundation models struggle to generalize as public 3D datasets are small and do not cover the broad diversity of medical procedures, conditions, anatomical regions, and imaging protocols. We address this by creating a representation learning method that instead anticipates strong domain shifts at training time itself. We first propose a data engine that synthesizes highly variable training samples that would enable generalization to new biomedical contexts. To then train a single 3D network for any voxel-level task, we develop a contrastive learning method that pretrains the network to be stable against nuisance imaging variation simulated by the data engine, a key inductive bias for generalization. This network’s features can be used as robust representations of input images for downstream tasks and its weights provide a strong, dataset-agnostic initialization for finetuning on new datasets. As a result, we set new standards across both multimodality registration and few-shot segmentation, a first for any 3D biomedical vision model, all without (pre-)training on any existing dataset of real images.
@@ 1 Introduction
Biomedical vision models trained on imaging studies with fixed protocols rarely generalize to new populations, medical procedures, and imaging devices. These domain shifts then necessitate practically infeasible reannotation and retraining cycles, especially for adaptation to new tasks. Further, volumetric annotated biomedical datasets are especially limited in sample size and focused on specific medical procedures, diseases, or scales of anatomy, leading current networks to overfit to a small subset of biomedical tasks. To overcome this data scarcity and heterogeneity, we present a representation learning framework driven by a synthetic data engine. Our approach yields a generalist 3D network that performs well on diverse voxel-level tasks across a range of unseen biomedical contexts in radiology.
@@
Current biomedical foundation models are trained by aggregating publicly available datasets to cover multiple domains (Butoi et al., 2023; Chen et al., 2024; Liu et al., 2023a; MH Nguyen et al., 2024; Pachitariu & Stringer, 2022; Xie et al., 2022). However, persistent gaps hinder their widespread adoption. For example, some methods operate only on specific modalities and regions that can be tractably scaled up in sample size, such as abdominal CT (Hamamci et al., 2024; Li et al., 2024a), and thus cannot learn general representations for most other domains, such as in utero fetal MRI. Others treat 2D slices of volumetric images as independent data points (Butoi et al., 2023; Ma & Wang, 2023), often constructing training sets with high inter-sample correlation and yielding models that fail to produce consistent 3D results. Further, existing foundation models almost exclusively focus on segmentation and classification and neglect other key vision tasks such as registration. To our knowledge, no biomedical vision foundation model has been demonstrated for multiple disparate 3D tasks yet.
@@
Contributions. This paper makes advances on two fronts. To gain robustness to large domain shifts in downstream deployment, we first propose a biomedically informed data engine whose samples encompass a wide range of appearances and semantics. This engine uses randomly sampled spatial configurations of biomedical shape templates to synthesize images with arbitrary resolutions, appearances, imaging physics, and crucially, minimal influence from any existing biomedical dataset. Unlike training on samples from GANs or diffusion models, which are limited to reproducing only their original training distribution, our engine synthesizes highly diverse samples useful for arbitrarily new biomedical contexts that we do not have training data for. We then develop a contrastive learning framework that uses paired samples from the data engine to pretrain a network for general voxel-level tasks in radiology using an inductive bias of approximate stability to nuisance imaging variation that does not change image semantics, a key property for generalization across datasets (Gruver et al., 2022).
@@
Our experiments demonstrate that the resulting features and weights enable broad generalization on the key biomedical tasks of 3D registration and segmentation across several diverse datasets. We achieve state-of-the-art unsupervised multimodality image registration by simply using the network’s approximately appearance invariant and pose equivariant representations (Fig. 1) to drive existing registration solvers. The proposed network can also be used as an off-the-shelf dataset-agnostic initialization for finetuning on any voxel-level task. Specifically, we demonstrate strong few-shot segmentation performance across several highly variable downstream datasets, thereby removing the need for the cumbersome dataset-specific self-supervised pretraining commonly used today. Our code and trained model weights are available at https://github.com/neel-dey/anatomix.
@@ 2 Related work
Figure 2: Data engine. A. We randomly sample binary labels as templates from a large database of anatomical segmentations to create 3D label ensembles of randomly deformed templates. B. Given a synthesized label ensemble and an appearance model, we synthesize two volumes to pretrain a network with a dense multi-view contrastive objective. C. Example synthetic training volumes produced by our data engine. These samples are not intended to be necessarily realistic, but rather to serve as diverse and useful training data for learning general tasks in arbitrary radiological domains.
@@
Domain randomization. To learn robustness to domain shifts at deployment, domain randomized generative models (Tobin et al., 2017) trade realism for diversity when generating training data for downstream models. For example, domain randomized methods for brain segmentation (Billot et al., 2023a; b; Gopinath et al., 2023; Hoopes et al., 2022b) and registration (Hoffmann et al., 2021; Iglesias, 2023) train on synthetic brains simulated from label maps and require large collections of expert brain annotations. Recent work in dataset-agnostic instance segmentation (Chollet et al., 2024; Dey et al., 2024) generalizes beyond brains by simulating both annotations and images using a pre-specified shape prior. We build on these concepts to develop a task-agnostic data engine that simulates images with highly variable appearances and physics using compositions of biomedical shape templates.
@@
Volumetric pretraining. Many methods pretrain on unannotated images from a dataset before supervised finetuning on a small labeled subset. These methods often employ self-supervised reconstructive (Chen et al., 2019a; Tang et al., 2022; Valanarasu et al., 2024; Zhou et al., 2023; 2021) and/or discriminative (Chaitanya et al., 2020; 2023; Ren et al., 2022; You et al., 2024) losses. However, these pretraining strategies exploit heuristics about their target datasets that often do not broadly generalize (Dong et al., 2021; Ren et al., 2022). Our approach instead yields a network that generalizes to arbitrary radiological datasets and does not require bespoke pretraining frameworks for each project. Lastly, recent biomedical foundation models trained on pooled datasets of 2D slices (Butoi et al., 2023; Ma & Wang, 2023; MH Nguyen et al., 2024; Wong et al., 2023) generally require interaction (via bounding boxes, scribbles, etc.), struggle with 3D consistency, and are restricted to segmentation. We instead directly train for general tasks using synthetic 3D volumes and do not work on interactive tasks.
@@ 3 Methods
Figure 3: Representation learning. Given a 3D label map and two corresponding synthetic volumes sampled from our data engine, we process them using a single UNet with shared weights. The UNet is pretrained contrastively at each decoder layer: for a randomly sampled anchor, features sampled from the same label in both volumes serve as positives, while features from other labels act as negatives.
@@
Data engine: label ensemble model (Fig. 2A). We create synthetic 3D label ensemble volumes by sampling from a repository of biomedical shape templates. As templates, we use the freely available $\sim$45,000 binary volumes from the TotalSegmentator dataset of 104 annotated organs in 1,204 CT volumes (Wasserthal et al., 2023). For each label ensemble volume, we iteratively populate a 3D volume with a random number of randomly sampled templates that are each then deformed and assigned a label corresponding to the sampling iteration. As radiological volumes are often surrounded by empty space (e.g., air), we simulate foreground and background separation by applying a binary foreground mask to two-thirds of the synthesized label ensembles by multiplying them with a randomly deformed binary sphere with a random radius and center. Further, as many radiological applications and domains have layer-like structures at the foreground-background interface (e.g., fat and skin), we randomly encase half of the foreground-masked volumes within envelope labels of random widths.
Data engine: appearance model (Fig. 2B). Given a 3D label ensemble $L$ with $K$ labels, we sample the intensities of two volumes $V_{1}$ and $V_{2}$ from two independent $K$-component Gaussian mixture models (GMMs) each with parameters $\{\mu_{k1},\sigma_{k1}^{2}\}_{k=1}^{K}$ and $\{\mu_{i2},\sigma_{i2}^{2}\}_{k=1}^{K}$, respectively, all of which are randomly drawn from uniform distributions. For each spatial index in $L$ with label $k$, we sample the initial intensities in $V_{1}$ and $V_{2}$ from $\mathcal{N}(\mu_{k1},\sigma_{k1}^{2})$ and $\mathcal{N}(\mu_{k2},\sigma_{k2}^{2})$, respectively. We then pointwise multiply them with Perlin-like noise (Perlin, 1985) to simulate spatial texture and augment using transformations relevant to biomedical volumes such as random bias fields, Fourier-spikes, Gamma shifts, blurring, Gibbs ringing, resolution degradations, noise, motion, flips, and affine warps. All intensity augmentations are sampled independently but the geometric augmentations are shared.
@@
Contrastive pretraining (Fig. 3). As our data engine provides exact label supervision, we develop a representation learning loss that is a spatial extension of multi-positive supervised contrastive learning (Khosla et al., 2020). To pretrain network $F:\mathbb{R}^{H\times W\times D}\rightarrow\mathbb{R}^{H\times W\times D\times C}$ where $H,W,$ and $D$ are spatial dimensions and $C$ is the number of output features, we use an inductive bias of voxel-level features within a 3D shape having similar spatial representations, regardless of differences in appearance. We assume that an anchor spatial index $i\in I$ (where $I=\{1,\dots,2HWD\}$, i.e., voxels pooled from $V_{1}$ and $V_{2}$) with features $f_{i}\in\mathbb{R}^{C}$ in label $k$ should have similar representations to all other indices in label $k$ in both $F(V_{1})$ and $F(V_{2})$ and dissimilar representations to indices from other labels. As in Chen et al. (2020a), we use a non-linear projection $Z:\mathbb{R}^{H\times W\times D\times C}\rightarrow\mathbb{R}^{H\times W\times D\times C_{Z}}$ on $F$’s outputs followed by an $L_{2}$-normalization when computing the contrastive loss,
@@
Figure 4: Multi-modality 3D registration. Using our representations with the ConvexAdam registration solver (“Ours”) yields accurate alignment of challenging pairs of intra-subject abdominal MRI-CT (top) and inter-subject cardiac MRI-CT (bottom) with large deformations across modalities.
@@
Downstream tasks: multi-modality registration. Gradient-based deformable registration objectives typically take the form of $\mathcal{L}_{reg}=d(V_{\mathrm{fixed}},V_{\mathrm{moving}}\circ\varphi)+\lambda\mathrm{Reg}(\varphi)$, where image $V_{\mathrm{moving}}$ is to be aligned to $V_{\mathrm{fixed}}$ by deformation $\varphi$, subject to regularization $\mathrm{Reg}(\cdot)$, and $d(\cdot)$ is an image dissimilarity score. To align images across imaging modalities, we simply replace the input volumes with our pretrained network’s features as in $\mathcal{L}_{reg}=d(F(V_{\mathrm{fixed}}),F(V_{\mathrm{moving}})\circ\varphi)+\lambda\mathrm{Reg}(\varphi)$. This approach is compatible with any existing high-performance registration solver, such as the ConvexAdam (Siebert et al., 2021; 2024) and ANTs (Tustison et al., 2020) frameworks used in our experiments below.
Downstream tasks: few-shot segmentation. For $N$-label few-shot semantic segmentation, we finetune the pretrained network $F$ on a small number of annotated volumes with an additional convolutional layer with softmax activation and $N$ output channels. We optimize the network using an equally weighted sum of the soft Dice and cross-entropy losses (Isensee et al., 2021; Taghanaki et al., 2019). To extract strong performance for all baselines in the challenging setting of finetuning on only one or few annotated volumes, we use extensively-tuned augmentation pipelines and finetune all layers of each network for a high number of iterations (37,500) with cosine learning rate decay.
@@ 4 Experiments
The pretrained network’s output features for inter and intra-subject volume pairs with semantically similar content subject to strong domain shifts are visualized in Fig. 1. Below, we present experiments that investigate the utility of our learned representations for multi-modality registration, the network weights as a pretrained initialization for few-shot segmentation, and analyze our modeling decisions.
@@ 4.1 Unsupervised multi-modality deformable 3D registration
Data and setup. We use the Learn2Reg AbdomenMRCT (Hering et al., 2021) (L2RAb) and MM-WHS (Gao et al., 2023; Zhuang, 2018; Zhuang et al., 2019) datasets to benchmark MRI to CT volume registration. L2RAb is an abdominal registration benchmarking dataset, whose publicly available portion provides eight affine-aligned intra-subject MRI and CT pairs of size $192\times 160\times 192$ at $2\times 2\times 2\mathrm{mm}^{3}$ resolution, with labels for four organs. MM-WHS, originally a heart segmentation dataset, contains 20 annotated MRIs and CTs (from distinct subjects) and we affine-align all volumes to a common space of size $160\times 160\times 128$ at $1.142\times 1.142\times 1.283\mathrm{mm}^{3}$ resolution (see App. B.4.4 for further details). The registration experiments are unsupervised and we split L2RAb and MM-WHS into 1/7 and 5/15 validation/testing pairs, respectively, where the validation pair(s) are only used for tuning registration hyperparameters.
Baselines and evaluation. We compare unsupervised and training-free multimodality registration frameworks. Our iterative baselines include the widely-used ANTs library (Avants et al., 2008; Tustison et al., 2020) with mutual information loss (Mattes et al., 2001; Wells III et al., 1996) (ANTs-MI) and the state-of-the-art (Hering et al., 2021) multimodality method, ConvexAdam (Siebert et al., 2021). We further use two dataset-agnostic registration networks, SynthMorph-shapes (Hoffmann et al., 2021) and uniGradICON (Tian et al., 2024a), with the latter also using optional instance optimization (uniGradICON+IO). For evaluation, we report post-registration volume overlap (Dice) of anatomical structures. We also assess deformation inverse consistency using the percentage of folding voxels, where $\mathrm{det}(J_{\varphi})<0$, for Jacobian $J_{\varphi}$ of the estimated warp $\varphi$. Folding percentages below 0.5% of all voxels are generally considered negligible and methods producing higher Dice values while staying under this threshold are preferred (Qiu et al., 2021).
@@
Results. Figs. 4 and 5 present results on held-out testing pairs. ANTS-Ours strongly improves upon the typically-used ANTS-MI, with 26 and 5 points of median Dice improvement on L2RAb and MM-WHS, respectively, while maintaining nearly identical low folding characteristics. Further, driven by our network features, ConvexAdam-Ours outperforms all methods in terms of volume overlap and improves on its base ConvexAdam method by 11 and 6 Dice points, with the same folding ratios. In contrast, the SynthMorph-shapes and uniGradICON+IO methods perform well on MM-WHS (where all hearts are roughly centered) but cannot handle the larger deformations in L2RAb. They do, however, yield nearly diffeomorphic transformations, producing almost zero folds. Lastly, without iterative optimization, uniGradICON demonstrates limited generalization across large intensity-based domain gaps. We note that all methods produce folding percentages that are under the threshold of 0.5% folding voxels. Additional grid search results on the validation sets are in App. A.2.